Solve each quadratic equation by completing the square.
4x(squared)=-2x+1
I get it where there are variables on one side and constants on the other, but the fractions are throwing me off.

Answer by stanbon(75887)

Solve each quadratic equation by completing the square.
4x(squared)=-2x+1
--
4x^2 + 2x = 1
4(x^2 + (1/2)x + ? = 1 + 4*?
---
The 3rd term is ((1/2)b)^2 = [(1/2)(1/2)]^2 = (1/4)^2
--------------------------------------------------
4(x^2 + (1/2)x + (1/4)^2) = 1 + 4(1/4)^2
4(x+(1/4)^2 = 1 + 4(1/16)
4(x+(1/4))^2 = 1 + (1/4)
4(x+(1/4))^2 = 5/4
(x+(1/4))^2 = 5/16
x+(1/4) = (1/4)sqrt(5)
x = -(1/4+-(1/4sqrt(5)
x = (-1/4)[1+-sqrt(5)]
============================
Cheers,
Stan H.